OpenStudy (ac3):

Change the cartesian integral into an equivalent polar integral. Then evaluate the polar integral.

OpenStudy (ac3):

\[\int\limits_{-1}^{1}\int\limits_{0}^{\sqrt{1-x ^{2}}}dydx\]

OpenStudy (ac3):

I got pi for my answer.

OpenStudy (ac3):

The correct answer is pi/2

OpenStudy (ac3):